OpenStudy (kitkat1):

Which of the following describes a linear function? It is V shaped and passes through the origin. It is a straight line in one portion and a curve in another portion. Its y-values decrease at a constant rate as its x-value increases. Its y-values increase at a nonconstant rate as its x-value increases.

OpenStudy (kitkat1):

Are u sure?

OpenStudy (kitkat1):

Does anyone else agree with this answer?

OpenStudy (anonymous):

third

OpenStudy (anonymous):

forth can't be because of non-constant

OpenStudy (kitkat1):

Okay

OpenStudy (anonymous):

medal me ok?

OpenStudy (kitkat1):

Who answers me second question will get the medal

OpenStudy (anonymous):

i will

OpenStudy (kitkat1):

Which set of ordered pairs represents a function?

OpenStudy (anonymous):

where is the function?

OpenStudy (kitkat1):

it doesnt let me send them wait a litle

OpenStudy (kitkat1):

{(0, 1), (1, 3), (1, 5), (2, 8)}

OpenStudy (anonymous):

the function is the ordered pair where no one y- value has more than two x ones

OpenStudy (kitkat1):

{(0, 0), (1, 2), (2, 6), (2, 8)}

OpenStudy (kitkat1):

{(0, 0), (0, 2), (2, 0), (2, 4)}

OpenStudy (kitkat1):

{(0, 2), (1, 4), (2, 6), (3, 6)

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

Plot them

OpenStudy (anonymous):

fortget

OpenStudy (kitkat1):

so the last one are u sure?

OpenStudy (anonymous):

you can determine that without plotting thousand times faster yes i'm sure

OpenStudy (kitkat1):

Can u help me with just one more or no?

OpenStudy (anonymous):

yes

OpenStudy (kitkat1):

If y – 3 = 3x, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs?

OpenStudy (anonymous):

what are all those question marks?

OpenStudy (kitkat1):

If y – 3 = 3x, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs

OpenStudy (kitkat1):

{(0, 3), (1, 6), (2, 9)}

OpenStudy (kitkat1):

{(3, 0), (6, 1), (9, 2)}

OpenStudy (kitkat1):

{(1, 3), (2, 6), (3, 9)}

OpenStudy (kitkat1):

{(3, 1), (6, 2), (9, 3)}

OpenStudy (kitkat1):

Do u see the question right now?

OpenStudy (anonymous):

the second one

OpenStudy (kitkat1):

if y - 3=3x which of the following sets represents possible inputs and outputs of the function represented as ordered pairs.

OpenStudy (kitkat1):

so the second one okay ill let u know if I get those questions right

OpenStudy (kitkat1):

How can I make u my fan I dont know how to do that

OpenStudy (anonymous):

click on ttb123456789 abd then click FaAN

OpenStudy (kitkat1):

I did thanks for help